# Theoretical analysis of photon detection mechanism in superconducting single-photon detectors

### v1

@@ I Introduction
The superconducting single-photon detector (SSPD) has many advantages, such as fast response, detection efficiency, low jitter, and low dark count [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11]. The extremely high photon detection capability is essential for photonic quantum information technologies. Nevertheless, the photon detection mechanism of the SSPD is not adequately understood. Therefore, currently there are no specific guidelines for the selection of superconducting materials and the design of devices to further improve performance. Typically, SSPD devices are composed of a superconducting nanostrip with a meander pattern (superconducting nanowire single-photon detector, SNSPD), and there are still restrictions regarding the polarization dependence and productivity. In this paper, our aim is to investigate the SSPD mechanism to overcome this situation.
@@
The SNSPD mechanism has been widely believed to be described by a hotspot model. According to this model, the device size for high-performance detection is severely limited, and for this reason the nanostrip structure is usually fabricated. However, recent experiments have shown that single-photon detection can also be performed with a micrometer-sized device [12, 13] (superconducting wide-strip photon detector, SWSPD). In particular, the authors in Ref. [13] propose a high critical current bank structure to suppress the intrinsic dark count caused by the nonuniform distribution of superconducting currents in the microstrip. The experimental successes for efficient photon detection by using the wide strip provide a good opportunity to reconsider the traditional hotspot model for the SSPD and SNSPD mechanisms.
@@
In the hotspot model, the superconductivity is locally suppressed by the incident photon whose energy scale is two or three orders of magnitude larger than the superconducting gap. In this case, without going into the details of the microscopic electronic states, it can be expected that the suppression is a thermal process. We refer to this process as initial hotspot formation. The superconducting current in the SSPD device flows outside the hotspot region. If the device is composed of a thin wire, the increase in the superconducting current density outside the hotspot region may exceed a critical value, leading to local breaking of the superconducting state. This process can be observed by the voltage between the two wire terminals. Although the formation of the initial hotspot region is plausible, the description of the subsequent processes seems to be insufficient to understand the detection mechanism in the micrometer-sized SSPD devices. In the subsequent processes, the hotspot model focuses on the expansion of the region, and there are many previous studies on the hotspot dynamics in terms of the heat-diffusion equation [14, 15, 16, 17]. However, analyzing the thermal distribution alone is not sufficient for a detailed understanding of electromagnetic response.
@@
An alternative possible mechanism is based on the formation of the vortex–antivortex pair. This idea was originally proposed in Refs. [18, 19], but to the best of our knowledge, no detailed study has been reported so far. In this vortex–antivortex model, the dynamics of vortices and antivortices perpendicular to the current flow direction result in a voltage change. A vortex-assisted photon-count mechanism was also proposed in Ref. [20], inspired by the single-vortex-crossing mechanism for dark counts [21]. A recent study [22] reports a unified theory of dark‐count rate and system‐detection efficiency based on a vortex-crossing framework. The photon detection mechanism based on this model is expected to be less sensitive to the finite-size limitations, consistent with the realization of SWSPD. The creation of vortex–antivortex pairs is also observed by the local heating effect of a scanning tunneling microscope tip [23]. Theoretical confirmation of pair generation would represent an important step toward developing a comprehensive understanding of the physics of SSPD devices. At the same time, we aim to clarify how the model parameters differ between the hotspot and vortex–antivortex models.
@@
The examination of the vortex–antivortex mechanism in terms of the TDGL equation is reported in Ref. [25]. The authors numerically solved the two-dimensional TDGL equation coupled with the Poisson equation for the electric potential and the heat-diffusion equation. They do not take into account the vector potential for the magnetic field. This assumption is valid in the large-$\kappa$ limit. They take the coherence length at zero temperature, the initial radius of the hotspot, and the width of the film (perpendicular to the current-flow direction) as $\xi_{0}=7.5$ nm, $R_{\mathrm{init}}=9$ nm, and $w=13\xi_{0}\sim 52\xi_{0}$, respectively. The local temperature increase is determined by energy conservation $E_{\mathrm{photon}}=\pi R_{\mathrm{init}}^{2}dC_{v}\Delta T$, where $C_{v}$ is a heat capacity, $d$ is a thickness of the sample, and $\Delta T$ is a temperature increase due to photon injection. At current larger than a threshold (the authors call this detecting current), they found that the superconducting state collapses starting from the appearance of a vortex–antivortex pair in the center of the initial hotspot region. Lorentz force causes their motion that heats the system locally and gives rise to a normal domain. The authors particularly focus on the dependence of detecting current on $\Delta T$ and $w$. In this study, $\xi_{0}$ is comparable to $R_{\mathrm{init}}$, but we consider that the initial hotspot size strongly affects the stability of the formation of a vortex–antivortex pair. Therefore, the parameters should be determined more carefully. Furthermore, the authors do not comment on why the traditional hotspot mechanism is not consistent with their analysis.
@@
Motivated by the present status mentioned above, we examine the dynamics of type-II superconductors with a bias current using the two-dimensional TDGL and the Maxwell equations. The photon injection is treated phenomenologically by the temperature increase inside the initial hotspot. The important parameter in this study is the initial hotspot radius by which the amount of photon energy is controlled. We find that certain parameter ranges can be explained by the hotspot model, while other parameter ranges are governed by the generation and annihilation of a vortex–antivortex pair. The photon detection is possible for an initial hotspot radius that exceeds a threshold value. We demonstrate that the pair generation occurs near the threshold. The flow of the pair perpendicular to the current direction finally leads to the formation of a normal region responsible for photon detection. During the formation process, the voltage change shows oscillation that originates from repeated penetration of vortices and antivortices from the sample edges. The voltage change for the GL parameter close to the transition point from type-II to type-I superconductor shows anomalous oscillations that are not associated with the vortex–antivortex pair. We also examine the effects of spatially non-uniform current density on the voltage change and the superconducting order parameter in order to understand the behavior of wide-strip single-photon detectors. The estimated values of incident photon energy and response time for photon detection are reasonable in comparison with experiments. The present comprehensive examination provides useful guidelines for flexible design of device structures.
@@ II Model
Figure 1: Illustration of our system. The red and blue areas correspond to superconductor and metallic electrodes, respectively. The system size and current-flow direction are indicated, respectively. The superconducting region is given by $0\leq x\leq L_{x}$ and $0\leq y\leq L_{y}$. The origin of the initial hotspot with radius $R_{\mathrm{init}}$ is set to the center of the superconductor.
@@
In this study, the photon injection is treated phenomenologically. A simple method is to consider heating in a fixed area due to the photon injection. We introduce the following temperature difference in and out of the initial hotspot:
@@
To preserve the gauge invariance of these equations even after discretization, we employ the link-variable method [26]. In the numerical simulation, the system is discretized into $N_{x}\times N_{y}$ meshes with a mesh spacing of $\delta=0.25\xi_{0}$ in both directions, i.e., $\delta=L_{x}/N_{x}=L_{y}/N_{y}$. For even integers $N_{x}$ and $N_{y}$, the numerical grid points are defined as $\bm{r}_{i_{x},i_{y}}=(i_{x}-1/2,i_{y}-1/2)\delta$, where $i_{\mu=x,y}=1,\cdots,N_{\mu}$. Note that there are no grid points located at $x=L_{x}/2$ or $y=L_{y}/2$ due to the above discretization; in the following calculations, these positions are given by $(L_{x}-\delta)/2$ and $(L_{y}-\delta)/2$, respectively; nevertheless, for simplicity, we denote them as $L_{x}/2$ and $L_{y}/2$, as long as no confusion arises. In particular, as shown below, such a shift in $\bm{r}_{c}$ leads to asymmetric dynamics. However, this asymmetry is not essential for photon detection, and we have confirmed that the symmetric dynamics are recovered when the hotspot center is set exactly at the system center, $\bm{r}_{c}=(L_{x}/2,L_{y}/2)$.
@@ III.1 Dependence of $R_{\mathrm{init}}$ on voltage change
Let us present numerical results. We first examine the dependence of $R_{\mathrm{init}}$ on the photon detection performance for $j_{\mathrm{bias}}/j_{0}=0.28$ and $\kappa=10$. The bias current density is set to be slightly lower than the critical current density (the critical value of $j_{\mathrm{bias}}/j_{0}$ is about $0.29$), where $j_{0}=(e^{\ast}/m^{\ast})\Delta_{\infty}^{2}(0)(\hbar/\xi_{0})$. The photon detection performance is characterized by the time evolution of the voltage change $\Delta V$, which is normalized by $V_{0}=H_{c2}\xi_{0}^{2}/t_{0}$. Here, we do not consider any thermal relaxation processes in which $\Delta V$ again goes to zero toward the next photon detection, since we particularly focus on the detection mechanism. We believe that our conclusions do not change so much even if we additionally introduce these thermal relaxation processes.
@@ III.2 Dynamics of amplitude and phase of superconducting order parameter
Figure 5: Time evolution of spacetime profile of $\left|\tilde{\Delta}\left(x=L_{x}/2,y\right)\right|$ for $150\leq t/t_{0}\leq 500$. The blight region corresponds to stable superconducting state. The position, $y=10$, corresponds to the hotspot center. (a) $R_{\mathrm{init}}/\xi_{0}=3.0$, $j_{\mathrm{bias}}/j_{0}=0.28$, and $\kappa=10$. (b) $R_{\mathrm{init}}/\xi_{0}=3.8$, $j_{\mathrm{bias}}/j_{0}=0.25$, and $\kappa=5$.
Let us examine the origin of the oscillation in terms of the order parameter. For this purpose, we first present the time evolution of $\left|\tilde{\Delta}\left(x=L_{x}/2,y\right)\right|$ (passing through the hotspot area) in Fig. 5(a), where we take $R_{\mathrm{init}}/\xi_{0}=3.0$, $j_{\mathrm{bias}}/j_{0}=2.8$, and $\kappa=10$. For $t/t_{0}\leq 250$, there are no significant changes outside the initial hotspot region, but the dark region begins to develop inside the hotspot. This dark region flows along the $y$ axis and arrives at the upper and lower edges of the superconductor around $t/t_{0}=250$. After these initial dynamics, the dark and blight areas appear repeatedly in the intermediate region between the sample edges and the initial hotspot region. The repetition period is equal to the oscillation period observed in Fig. 2.
For comparison with the case where $\kappa$ is smaller, we present $\left|\tilde{\Delta}\left(x=L_{x}/2,y\right)\right|$ for $j_{\mathrm{bias}}/j_{0}=0.25$, $\kappa=5$, and $R_{\mathrm{init}}/\xi_{0}=3.8$ in Fig. 5(b). The oscillatory behavior of the dark and blight areas is also observed. However, before the flow of the dark area in the initial hotspot toward the edges, another dark region begins to penetrate from the lower edge of the superconductor around $t/t_{0}=350$. After that, these dark areas merge with each other. The dynamics of the (anti)vortex penetration is spatially asymmetric about $y=L_{y}/2$. This is due to the slight shift of the initial hotspot from the center of the system in the lattice discretization as mentioned at the end of Sect. II. We have confirmed numerically that this asymmetry does not occur when the hotspot is exactly located at the center of the system. As shown in the next paragraph, the dark regions originates in the vortex–antivortex pair generated in the initial hotspot for $t/t_{0}\lesssim 250$ and the penetrating vortices and antivortices for $t/t_{0}\gtrsim 250$.
@@
To visualize the generation of the vortex–antivortex pair in the initial hotspot region and the formation of the normal region, we present color maps of the amplitude of the superconducting order parameter $\tilde{\Delta}(x,y)$ for various times in Fig. 6(a). We take $R_{\mathrm{init}}/\xi_{0}=3.0$, $j_{\mathrm{bias}}/j_{0}=0.28$, and $\kappa=10$. We find that the normal region is finally formed across the upper and lower ends of the sample, and this is consistent with the voltage change in Fig. 3. During this process, we observe two small dark spots generated in the initial hotspot and moving separately toward the ends of the sample [see panels (ii)–(iv) of Fig. 6(a)]. These spots are smaller than $R_{\mathrm{init}}/\xi_{0}=3$, and are comparable in size to the vortex core (of order $\xi_{0}$). After the dark spots reach the sample edges around $t/t_{0}=250$, the solid normal region is finally formed across the upper and lower edges [see Fig. 6(a)(v)]. Our observation cannot be simply understood in terms of the hotspot model.
We also plot the phase of the order parameter in Fig. 6(b). We clearly observe that the two small dark spots correspond to quantum vortices with opposite phase windings ($\pm 2\pi$ phase rotations) [see panels (ii)–(iv) of Fig. 6(b)]. Thus, we conclude that they are a vortex–antivortex pair. We find that the intermediate region between the two singularities expands as $t$ increases. At $t/t_{0}=250$ [see Fig. 6(b)(iv)], this region is almost in contact with the lower edge of the sample. According to Fig. 2, a sharp increase in $\Delta V$ occurs around this time. We cannot clearly observe the presence of penetrating vortices and antivortices at least up to $t/t_{0}\leq 250$. After $t/t_{0}=250$, $\Delta V$ starts oscillating, as shown in Fig. 2.
By careful analysis of the time evolution of the order parameter, we have confirmed repeated penetration of vortex-antivortex pairs from the outside of the superconductor, and the correspondence between the period of the oscillation in $\Delta V$ and that of the repeated penetration. For instance, the first local minimum and the second local maximum in $\Delta V$ are located around $t/t_{0}=254.5$ and $t/t_{0}=261$, respectively, and the vortex and antivortex penetrate from the outside around $t/t_{0}=258$. The time when $\Delta V$ reaches its local minimum corresponds to the onset of the decrease in the edge order parameter, serving as a precursor to the penetration of a vortex or an antivortex, while its local maximum is reached after the penetrating vortex and antivortex annihilate near the center.

### v2

@@ I Introduction
The superconducting single-photon detector (SSPD) has many advantages, such as fast response, detection efficiency, low jitter, and low dark count [Goltsman, Goltsman2, Yamashita, Hadfield, Miki, Miki2, Yamashita2, Yamashita3, Yamashita4, Miki3, Yamashita5]. The extremely high photon detection capability is essential for photonic quantum information technologies. Nevertheless, the photon detection mechanism of the SSPD is not adequately understood. Therefore, currently there are no specific guidelines for the selection of superconducting materials and the design of devices to further improve performance. Typically, SSPD devices are composed of a superconducting nanostrip with a meander pattern (superconducting nanowire single-photon detector, SNSPD), and there are still restrictions regarding the polarization dependence and productivity. In this paper, our aim is to investigate the SSPD mechanism to overcome this situation.
@@
The SNSPD mechanism has been widely believed to be described by a hotspot model. According to this model, the device size for high-performance detection is severely limited, and for this reason the nanostrip structure is usually fabricated. However, recent experiments have shown that single-photon detection can also be performed with a micrometer-sized device [Korneeva, Yabuno] (superconducting wide-strip photon detector, SWSPD). In particular, the authors in Ref. [Yabuno] propose a high critical current bank structure to suppress the intrinsic dark count caused by the nonuniform distribution of superconducting currents in the microstrip. The experimental successes for efficient photon detection by using the wide strip provide a good opportunity to reconsider the traditional hotspot model for the SSPD and SNSPD mechanisms.
@@
In the hotspot model, the superconductivity is locally suppressed by the incident photon whose energy scale is two or three orders of magnitude larger than the superconducting gap. In this case, without going into the details of the microscopic electronic states, it can be expected that the suppression is a thermal process. We refer to this process as initial hotspot formation. The superconducting current in the SSPD device flows outside the hotspot region. If the device is composed of a thin wire, the increase in the superconducting current density outside the hotspot region may exceed a critical value, leading to local breaking of the superconducting state. This process can be observed by the voltage between the two wire terminals. Although the formation of the initial hotspot region is plausible, the description of the subsequent processes seems to be insufficient to understand the detection mechanism in the micrometer-sized SSPD devices. In the subsequent processes, the hotspot model focuses on the spatiotemporal density change of quasiparticles or the expansion of the region. There are previous studies on the hotspot dynamics using the kinetic-equation approach or the heat-diffusion equation [Kadin3, Semenov, Semenov2, Engel, Kozorezov, Marsili, Vodolazov]. However, analyzing the thermal distribution alone is not sufficient for a detailed understanding of electromagnetic response.
@@
An alternative possible mechanism is based on the formation of the vortex–antivortex pair. This idea was originally proposed in Refs. [Kadin, Kadin2], but to the best of our knowledge, no detailed study that directly visualizes their dynamics has been reported so far. In this vortex–antivortex model, the dynamics of vortices and antivortices perpendicular to the current flow direction result in a voltage change. Vortex-assisted mechanisms were also proposed in Refs. [Qiu, Bartolf, Hofherr, Bulaevskii2, Bulaevskii, Zotova2, Vodolazov2, Engel2, Saman], inspired by the single-vortex-crossing mechanism for dark count [Hofherr, Bulaevskii2]. A recent study [He] reports a unified theory of dark‐count rate and system‐detection efficiency based on a vortex-crossing framework. The photon detection mechanism based on this model is expected to be less sensitive to the finite-size limitations, consistent with the realization of SWSPD. The creation of vortex–antivortex pairs is also observed by the local heating effect of a scanning tunneling microscope tip [Ge]. Theoretical confirmation of pair generation would represent an important step toward developing a comprehensive understanding of the physics of SSPD devices. At the same time, we aim to clarify how the model parameters differ between the hotspot and vortex–antivortex models.
@@
The examination of the vortex–antivortex mechanism in terms of the TDGL equation is reported in Ref. [Zotova]. The authors numerically solved the two-dimensional TDGL equation coupled with the Poisson equation for the electric potential and the heat-diffusion equation. They do not take into account the vector potential for the magnetic field. This assumption is valid in the large-$\kappa$ limit. They take the coherence length at zero temperature, the initial radius of the hotspot, and the width of the film (perpendicular to the current-flow direction) as $\xi_{0}=7.5$ nm, $R_{\mathrm{init}}=9$ nm, and $w=13\xi_{0}\sim 52\xi_{0}$, respectively. The local temperature increase is determined by energy conservation $E_{\mathrm{photon}}=\pi R_{\mathrm{init}}^{2}dC_{v}\Delta T$, where $C_{v}$ is a heat capacity, $d$ is a thickness of the sample, and $\Delta T$ is a temperature increase due to photon injection. At current larger than a threshold (the authors call this detecting current), they found that the superconducting state collapses starting from the appearance of a vortex–antivortex pair in the center of the initial hotspot region. Lorentz force causes their motion that heats the system locally and gives rise to a normal domain. The authors particularly focus on the dependence of detecting current on $\Delta T$ and $w$. In this study, $\xi_{0}$ is comparable to $R_{\mathrm{init}}$, but we consider that the initial hotspot size strongly affects the stability of the formation of a vortex–antivortex pair. Therefore, the parameters should be determined more carefully. Furthermore, the authors do not comment on why the traditional hotspot mechanism is not consistent with their analysis.
@@
Motivated by the present status mentioned above, we examine the dynamics of type-II superconductors with a bias current using the two-dimensional TDGL and the Maxwell equations. The photon injection is treated phenomenologically by the temperature increase inside the initial hotspot region. The important parameter in this study is the initial hotspot radius by which the amount of photon energy is controlled. We find that certain parameter ranges can be explained by the hotspot model, while other parameter ranges are governed by the generation and annihilation of a vortex–antivortex pair. The photon detection is possible for an initial hotspot radius that exceeds a threshold value. We demonstrate that the pair generation occurs near the threshold. The flow of the pair perpendicular to the current direction finally leads to the formation of a normal region responsible for photon detection. During the formation process, the voltage change shows oscillation that originates from repeated penetration of vortices and antivortices from the sample edges. The voltage change for the GL parameter close to the transition point from type-II to type-I superconductor shows anomalous oscillations that are not associated with the vortex–antivortex pair. We also examine the effects of spatially non-uniform current density on the voltage change and the superconducting order parameter in order to understand the behavior of wide-strip single-photon detectors. The estimated values of incident photon energy and response time for photon detection are reasonable in comparison with experiments. The present comprehensive examination provides useful guidelines for flexible design of device structures.
@@ II Model
Figure 1: Illustration of our system. The red and blue areas correspond to superconductor and metallic electrodes, respectively. The system size and current-flow direction are indicated, respectively. The superconducting region is given by $0\leq x\leq L_{x}$ and $0\leq y\leq L_{y}$. The center of the initial hotspot with radius $R_{\mathrm{init}}$ is set to the center of the superconductor.
@@
In this study, the photon injection is treated phenomenologically. A simple method is to consider heating in a fixed area due to the photon injection. We introduce the following temperature difference between the inside and outside of the initial hotspot:
@@
To preserve the gauge invariance of these equations even after discretization, we employ the link-variable method [Kato]. In the numerical simulation, the system is discretized into $N_{x}\times N_{y}$ meshes with a mesh spacing of $\delta=0.25\xi_{0}$ in both directions, i.e., $\delta=L_{x}/N_{x}=L_{y}/N_{y}$. For even integers $N_{x}$ and $N_{y}$, the numerical grid points are defined as $\bm{r}_{i_{x},i_{y}}=(i_{x}-1/2,i_{y}-1/2)\delta$, where $i_{\mu=x,y}=1,\cdots,N_{\mu}$. Note that there are no grid points located at $x=L_{x}/2$ or $y=L_{y}/2$ due to the above discretization; in the following calculations, these positions are given by $(L_{x}-\delta)/2$ and $(L_{y}-\delta)/2$, respectively; nevertheless, for simplicity, we denote them as $L_{x}/2$ and $L_{y}/2$, as long as no confusion arises. In particular, as shown below, such a shift in $\bm{r}_{c}$ leads to asymmetric dynamics. However, this asymmetry is not essential for photon detection, and we have confirmed that the symmetric dynamics are recovered when the hotspot center is set exactly at the system center, $\bm{r}_{c}=(L_{x}/2,L_{y}/2)$.
@@ III.1 Dependence of $R_{\mathrm{init}}$ on voltage change
Let us present numerical results. We first examine the dependence of $R_{\mathrm{init}}$ on the photon detection performance for $j_{\mathrm{bias}}/j_{0}=0.28$, $j_{0}=(e^{\ast}\hbar/m^{\ast}\xi_{0})\Delta_{\infty}^{2}(0)$, and $\kappa=10$. The bias current density $j_{\mathrm{bias}}$ is set to be slightly lower than the critical current density $j_{c}$, where $0.30<j_{c}/j_{0}<0.31$. The photon detection performance is characterized by the time evolution of the voltage change $\Delta V$, which is normalized by $V_{0}=H_{c2}\xi_{0}^{2}/t_{0}$. Here, we do not consider any thermal relaxation processes in which $\Delta V$ again goes to zero toward the next photon detection, since we particularly focus on the detection mechanism. We believe that our conclusions do not change so much even if we additionally introduce these thermal relaxation processes.
@@ III.2 Dynamics of amplitude and phase of superconducting order parameter
Figure 5: Time evolution of spacetime profile of $\left|\tilde{\Delta}\left(x=L_{x}/2,y\right)\right|$ [(a) upper panel] and that of $\Delta V$ [(b) lower panel] for $R_{\mathrm{init}}/\xi_{0}=3.0$, $j_{\mathrm{bias}}/j_{0}=0.28$, $\kappa=10$, and $200\leq t/t_{0}\leq 350$. Panel (b) is a magnified view of Fig. 2 for $R_{\mathrm{init}}/\xi_{0}=3.0$. The blight region in panel (a) corresponds to stable superconducting state. The position, $y/\xi_{0}=10$, corresponds to the hotspot center.
Let us examine the origin of the oscillation in terms of the order parameter. For this purpose, we first present the time evolution of $\left|\tilde{\Delta}\left(x=L_{x}/2,y\right)\right|$ (passing through the hotspot area) in Fig. 5(a), where we take $R_{\mathrm{init}}/\xi_{0}=3.0$, $j_{\mathrm{bias}}/j_{0}=2.8$, and $\kappa=10$. For $t/t_{0}\leq 250$, there are no significant changes outside the initial hotspot region, but a pair of localized dark spots begins to develop inside the hotspot. They are also localized in the $x$ direction. These dark spots propagate along the $y$ axis and reach the upper and lower edges of the superconductor around $t/t_{0}=250$. After $t/t_{0}=250$, two dark spots penetrate from the lower and upper edges of the superconductor and propagate toward each other, eventually merging near the center of the system. These processes occur repeatedly. Note that the dynamics is spatially asymmetric about $y=L_{y}/2$. This is due to the slight shift of the initial hotspot center relative to the system center, introduced by the lattice discretization as mentioned at the end of Sect. II. We have confirmed numerically that this asymmetry does not occur when the hotspot region is located such that its center coincides with the center of the system.
For comparison, we also plot $\Delta V$ in the same time domain as Fig. 5(b). Around $t/t_{0}=250$, $\Delta V$ shows its first local maximum. At the same time, the dark spots generated within the initial hotspot region reach the sample edges, as shown in Fig. 5(a). The other local maxima in $\Delta V$ appear, when the pairs of the penetrating dark spots in Fig. 5(a) merge with each other near the center of the sample. The penetration of each pair occurs between times when $\Delta V$ reaches its local minimum and its subsequent local maximum. As shown in the next paragraph, the dark spots originates in the vortex–antivortex pair generated in the initial hotspot for $t/t_{0}\lesssim 250$ and in the penetrating vortices and antivortices for $t/t_{0}\gtrsim 250$.
@@
To visualize the generation of the vortex–antivortex pair in the initial hotspot region and the subsequent formation of the normal region, we present in Fig. 6(a) color maps of the amplitude of the superconducting order parameter $\tilde{\Delta}(x,y)$ at several representative times, obtained for $R_{\mathrm{init}}/\xi_{0}=3.0$, $j_{\mathrm{bias}}/j_{0}=0.28$, and $\kappa=10$. We find that the normal region is finally formed across the upper and lower ends of the sample, and this is consistent with the voltage change in Fig. 3. During this process, we observe two small dark spots generated in the initial hotspot and moving separately toward the ends of the sample [see panels (ii)–(iv) of Fig. 6(a)]. The radii of these spots are smaller than $R_{\mathrm{init}}/\xi_{0}=3.0$, and are comparable in size to the vortex core (of order $\xi_{0}$). After the dark spots reach the sample edges around $t/t_{0}=250$, the solid normal region is finally formed across the upper and lower edges [see Fig. 6(a)(v)]. Our observation cannot be simply understood in terms of the hotspot model.
We also plot the corresponding phases of the order parameter in Fig. 6(b). We clearly observe that the two small dark spots correspond to quantum vortices with opposite phase windings ($\pm 2\pi$ phase rotations) [see panels (ii)–(iv) of Fig. 6(b)]. Thus, we conclude that they are a vortex–antivortex pair. We find that the intermediate region between the two singularities expands as $t$ increases. At $t/t_{0}=250$ [see Fig. 6(b)(iv)], this region is almost in contact with the lower edge of the sample. According to Fig. 2 or Fig. 5(b), a sharp increase in $\Delta V$ occurs and $\Delta V$ takes its first local maximum around this time. We cannot clearly observe the presence of penetrating vortices and antivortices at least up to $t/t_{0}\leq 250$. After $t/t_{0}=250$, $\Delta V$ starts oscillating, as shown in Fig. 2.
By careful analysis of the time evolution of the order parameter after $t/t_{0}=250$ (not shown here), we have confirmed repeated penetration of vortex-antivortex pairs from the outside of the superconductor, and this is the origin of the penetrating dark spots in Fig. 5(a). For instance, the first local minimum and the second local maximum in $\Delta V$ are located around $t/t_{0}=254.5$ and $t/t_{0}=261$, respectively, and the vortex and antivortex penetrate from the outside around $t/t_{0}=258$. The time when $\Delta V$ reaches its local minimum corresponds to the onset of the decrease in the edge order parameter, serving as a precursor to the penetration of a vortex or an antivortex, while its local maximum is reached after the penetrating vortex and antivortex annihilate near the center.